# Twinning Complex Networked Systems: Data-Driven Calibration of the mABCD Synthetic Graph Generator

### v1

@@ 4 Experiments
To test the proposed configuration retrieval method, two experiments using the goodness-of-fit framework proposed in Sec. 3 were performed. In both experiments, a real-world multilayer network, namely Freebase [22], was employed. The network consists of $\ell=3$ layers and includes 3,492 actors. The numbers of nodes in the consecutive layers are 3,479, 2,091, and 1,865, respectively. The corresponding numbers of edges per layer are 129,097, 7,099, and 5,948. The mean degree is 75.41, and the average closeness centrality is 0.1.
@@ 4.1 Experiment 1 – Dimensionality of the Reference Layer d
The objective of the first experiment was to evaluate the influence of the dimension of the reference layer ($d$) on the divergence scores between the original network and its produced twins. As such, the estimation procedure was performed independently for four different values of $d$, i.e. $d\in\{1,2,4,8\}$. Note that it directly affects the Bayesian optimisation component of the proposed method that is used to estimate $r$.
Figure 3: Divergence scores (log scale) for configuration retrieval with fixed $d=2^{k}$ dimension, $k\in\{0,1,2,3\}$; estimation with tuned $r$ and loss $\mathcal{D}_{r}(\hat{G},G)$.
Fig. 3 presents a comparison of the average values of the corresponding divergence metrics across the tested $d$. The scores associated with $\gamma$, $\beta$, $\xi$, and $\tau$ remain virtually unchanged, as expected, since these parameters were fixed during the optimisation of $r$. Their stability only confirms the reproducibility of the proposed approach. In contrast, a pronounced effect of $d$ is observed for the AMI-based interlayer correlation ($r$), resulting in non-monotonic divergence values with respect to increasing $d$, which is consistent with the hypothesis of parameter interdependence. Nevertheless, the lowest divergence is attained for $d=2$, which coincides with the design choice recommended by the authors of mABCD.
An additional, perhaps less anticipated, effect of a high $d$ value is the increased divergence in edge correlation ($R$). The increasing misalignment of communities between layers, when comparing the original network to the digital twin, creates unfavourable conditions for efficient link rewiring (phase 6 of the mABCD process). It should also be noted that relatively high values of $\gamma$ and $\beta$ can stem from the non necessarily power-law-like distribution of degrees and communities in Freebase. A potential improvement for that could be injecting sequences of them directly into mABCD and, by that, to pass over some of its phases.
@@ 5 Conclusions
The main contribution of this study is the proposal of a method for creating statistically faithful network twins from empirical multilayer datasets. We also introduce several divergence measures for assessing the quality of parameter estimation. The method is evaluated on a real-world network representing cooperation within the movie industry. The results provide preliminary evidence supporting the validity of the proposed approach; however, a key finding is that independent parameter estimation techniques encounter intrinsic accuracy limits, as the configuration parameters collectively determine the structural properties of networks generated by mABCD. This observation naturally motivates the development of joint prediction methods as a promising direction for future work.

### v2

@@ 4 Experiments
To test the proposed configuration retrieval method, two experiments using the framework proposed in Sec. 3 were performed. In both of them a real-world multilayer network Freebase [23], representing cooperation in the movie industry, was employed. The Freebase network captures relations between movies and people involved in their production (Actors, Writers, Directors). No explicit inter-layer edges are defined (which aligns with Def. 1); layers correspond to meta-path relations (MAM, MDM, MWM) over the same movie nodes. The dataset is used here only as an illustrative example, and its semantics are not considered further. The network has 3,492 actors; the numbers of nodes per layer are 3,479, 2,091, and 1,865, with 129,097, 7,099, and 5,948 edges, respectively. The mean degree is 75.41 and the average closeness centrality is 0.1.
@@ 4.1 Experiment 1 – Dimensionality of the Reference Layer d
The objective of the first experiment was to evaluate the influence of the dimension of the reference layer ($d$) on the divergence scores between the original network and its produced twins. As such, the estimation procedure was performed independently for four fixed values of $d$, i.e. $d\in\{1,2,4,8\}$. Note that it directly affects the Bayesian optimisation component of the proposed method that is used to estimate $r$.
Figure 3: Divergence scores for configuration retrieval with fixed $d=2^{k}$ dimension, $k\in\{0,1,2,3\}$; estimation with tuned $r$ and the optimisation loss $\mathcal{D}_{r}(\hat{G},G)$.
Fig. 3 presents a comparison of the average values of the corresponding divergence metrics across the tested $d$. The scores associated with $\gamma$, $\beta$, $\xi$, and $\tau$ remain virtually unchanged, as expected, since these parameters were fixed during the optimisation of $r$. Their stability only confirms the reproducibility of the proposed approach. In contrast, a pronounced effect of $d$ is observed for the AMI-based interlayer correlation ($r$), resulting in non-monotonic divergence values with respect to $d$, which is consistent with our hypothesis of parameters’ interdependence. Nevertheless, the lowest divergence is attained for $d=2$, which coincides with the design choice recommended by the authors of the mABCD model.
An additional, perhaps less anticipated, effect of using a high $d$ value is the increased divergence in edge correlation ($R$). The growing misalignment of communities between layers, when comparing the original network to the digital twin, creates unfavourable conditions for efficient link rewiring (phase 6 of the mABCD process). It should also be noted that relatively high values of $\gamma$ and $\beta$ can stem from the non necessarily power-law-like distribution of degrees and communities in Freebase. A potential improvement for that could be injecting sequences of them directly into mABCD and, by that, to pass over some of its phases.
@@ 5 Conclusions
The main contribution of this study is the proposal of a statistical network inference method from empirical multilayer datasets. We also introduce several divergence measures for assessing the quality of parameter estimation. The method is evaluated on a real-world network representing cooperation within the movie industry. The results provide preliminary evidence supporting the validity of the proposed approach; however, a key finding is that independent parameter estimation techniques encounter intrinsic accuracy limits, as the configuration parameters collectively determine the structural properties of networks generated by mABCD. This observation naturally motivates the development of joint prediction methods as a promising future direction.